OpenStudy (anonymous):

find the slope of the line that contains points (5,-1) and (-5,5) . write the equation in slope-intercept form.

OpenStudy (anonymous):

The slope of the line is given by \[slope=m=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\]The equation of the line is very similar:\[m=\frac{y-y_1}{x-x_1}\], which should be rearranged into slope-intercept form\[y=mx+c\]

OpenStudy (anonymous):

so the equation would be y-(-1)=(-3/5)(x-5) y2=-3/5x +3 What did I do wrong ???

OpenStudy (anonymous):

You had it correct up to here: y-(-1)=(-3/5)(x-5), so well done for that! \[y-(-1)=-\frac{3}{5}(x-5)\]\[\Leftrightarrow y+1=-\frac{3}{5}x+3\]\[\Leftrightarrow y=-\frac{3}{5}x-2\]

OpenStudy (anonymous):

so how does the y-(-1) become y+1?

OpenStudy (anonymous):

that's where i get confused,because y-(-1) oh, i guess i was multiplying 1(-1) but should be opposite of -1?

OpenStudy (anonymous):

\(y-(-1)\) means \(y-1(-1)\). \(-1\times-1=+1\Rightarrow y-(-1)=y+1\) \[\]